The capacitors in each circuit are fully charged b

The capacitors in each circuit are fully charged before the switch is closed. Rank, from longest to shortest, the length of time the bulbs (resistors) stay lit in each circuit.
Answer #1
Concepts and reason

The concept used to solve this problem is circuit analysis using the simplification of resistors and capacitors.

Initially, calculate the time constant for each of the circuits in figure A, figure B, figure C, figure D and figure E. Finally, rank the length of time the bulbs stay lit from longest to shortest by using the value of time constant for each circuit.

Fundamentals

The expression to calculate equivalent resistance for series combination of resistor is,

Rs=R1+R2{R_s} = {R_1} + {R_2}

Here, Rs{R_s} is the equivalent resistance in series and R1{R_1} , R2{R_2} are the resistances of individual resistors connected in series.

The expression to calculate equivalent resistance for parallel combination of resistor is,

Rp=R1R2R1+R2{R_p} = \frac{{{R_1}{R_2}}}{{{R_1} + {R_2}}}

Here, Rp{R_p} is the equivalent resistor in parallel.

The expression to calculate equivalent capacitance for parallel combination of capacitors is,

Cp=C1+C2{C_p} = {C_1} + {C_2}

Here, Cp{C_p} is the equivalent capacitance in parallel and C1{C_1} , C2{C_2} are the capacitances of individual capacitors connected in parallel.

The expression to calculate equivalent capacitance for series combination of capacitor is,

Cs=C1C2C1+C2{C_s} = \frac{{{C_1}{C_2}}}{{{C_1} + {C_2}}}

Here, Cs{C_s} is the equivalent capacitance of capacitors connected in series.

The expression for time constant in RC,

τ=RC\tau = RC

Here, τ\tau is the time constant, R is the resistance, and C is the capacitance.

The figure A has one resistance and one capacitor connected in series.

The expression for time constant in RC for figure A is,

τ=RC\tau = RC

The circuit has two capacitors in series with one resistor.

The expression for series combination of capacitor is,

Cs=C1C2C1+C2{C_s} = \frac{{{C_1}{C_2}}}{{{C_1} + {C_2}}}

Substitute CC for C1{C_1} and C2{C_2} to find Cs{C_s} .

Cs=CCC+C=C22C=C2\begin{array}{c}\\{C_s} = \frac{{CC}}{{C + C}}\\\\ = \frac{{{C^2}}}{{2C}}\\\\ = \frac{C}{2}\\\end{array}

The expression for time constant is,

τ=RC\tau = RC

Substitute C/2{\rm{C/2}} for C to find τ\tau .

τ=R(C2)=12RC\begin{array}{c}\\\tau = R\left( {\frac{C}{2}} \right)\\\\ = \frac{1}{2}RC\\\end{array}

The figure C has two capacitors connected in parallel with one resistance in series.

The expression for parallel combination of capacitor is,

Cp=C1+C2{C_p} = {C_1} + {C_2}

Substitute CC for C1{C_1} and C2{C_2} to find Cp{C_p} .

CP=C+C=2C\begin{array}{c}\\{C_P} = C + C\\\\ = 2C\\\end{array}

The expression for time constant is,

τ=RC\tau = RC

Substitute 2C2{\rm{C}} for C to find τ\tau .

τ=R(2C)=2RC\begin{array}{c}\\\tau = R\left( {2C} \right)\\\\ = 2RC\\\end{array}

The circuit has two capacitors in series with two resistances in parallel.

The expression for series combination of capacitor is,

Cs=C1C2C1+C2{C_s} = \frac{{{C_1}{C_2}}}{{{C_1} + {C_2}}}

Substitute CC for C1{C_1} C2{C_2} to find Cs{C_s} .

Cs=CCC+C=C22C=C2\begin{array}{c}\\{C_s} = \frac{{CC}}{{C + C}}\\\\ = \frac{{{C^2}}}{{2C}}\\\\ = \frac{C}{2}\\\end{array}

The circuit has two capacitors in series with one resistance in parallel.

The expression for parallel combination of resistor is,

Rp=R1R2R1+R2{R_p} = \frac{{{R_1}{R_2}}}{{{R_1} + {R_2}}}

Substitute R for R1{R_1} , R2{R_2} to find Rp{R_p} .

Rp=RRR+R=R22R=R2\begin{array}{c}\\{R_p} = \frac{{RR}}{{R + R}}\\\\ = \frac{{{R^2}}}{{2R}}\\\\ = \frac{R}{2}\\\end{array}

The expression for time constant is,

τ=RC\tau = RC

Substitute C/2{\rm{C/2}} for C and R/2R/2 for R to find τ\tau .

τ=R2(C2)=RC4\begin{array}{c}\\\tau = \frac{R}{2}\left( {\frac{C}{2}} \right)\\\\ = \frac{{RC}}{4}\\\end{array}

The circuit has two capacitors in parallel and two resistors in parallel.

The expression for parallel combination of capacitor is,

Cp=C1+C2{C_p} = {C_1} + {C_2}

Substitute CC for C1{C_1} , C2{C_2} to find Cp{C_p} .

CP=C+C=2C\begin{array}{c}\\{C_P} = C + C\\\\ = 2C\\\end{array}

The expression for parallel combination of resistor is,

Rp=R1R2R1+R2{R_p} = \frac{{{R_1}{R_2}}}{{{R_1} + {R_2}}}

Here, Rp{R_p} is the equivalent resistor in parallel and R1{R_1} R2{R_2} is the resistance.

Substitute R for R1{R_1} and R2{R_2} to find Rp{R_p} .

Rs=RRR+C=R22R=R2\begin{array}{c}\\{R_s} = \frac{{RR}}{{R + C}}\\\\ = \frac{{{R^2}}}{{2R}}\\\\ = \frac{R}{2}\\\end{array}

The expression for time constant is,

τ=RC\tau = RC

Substitute 2C{\rm{2C}} for C and R/2R/2 for R and solve for τ\tau .

τ=2R(C2)=RC\begin{array}{c}\\\tau = 2R\left( {\frac{C}{2}} \right)\\\\ = RC\\\end{array}

Ans:

Rank of the time constant of the circuits is C>A=E>B>DC > A = E > B > D